OpenStudy (anonymous):

What is the fraction 36/80 in lowest terms? WIll medal and fan

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

9/20. Do you need an explanation?

OpenStudy (anonymous):

na you don't have to

OpenStudy (welshfella):

you need to find the GCF of 36 and 80. That is the highest number which divides into 36 and 80. then divide top and bottom of the fraction by this GCF.

jabez177 (jabez177):

First you need to find the Lowest common denominator. Basically what number can both go into 36 and 80.

OpenStudy (anonymous):

Alright

jabez177 (jabez177):

Ah, the greatest common factor, sorry. :P I'm a bit rusty.

OpenStudy (anonymous):

Np

OpenStudy (welshfella):

@allenismyname - we should give an explanation anyway. See Code of Conduct

jabez177 (jabez177):

@Allenismyname, take a look at the code of conduct. It says we're not supposed to give out direct answers. We're supposed to guide them to the answer. Just a heads up. :P :)

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Ok. Divide the numerator and denominator by the GCF which is 4. 36/4=9 and 80/4=20= 9/20